# Magnetosphere Evolution and Precursor-Driven Electromagnetic Signals in Merging Binary Neutron Stars

### v1

@@ 3.1.5 Scaling Laws for Poynting Luminosity and the Binary Braking Index
Empirically fitting the gross trend from our simulations $\langle L\rangle\approx~{}\tilde{L}\Omega^{p}$, we find an expected correlation between $\lambda$ and the overall normalization $\tilde{L}$ of the Poynting luminosity. Namely, $\tilde{L}/(10^{42}\,B_{12}^{2}r_{12\,\rm km}^{2}\rm\,erg\,s^{-1})~{}\sim~{}0.8\exp{(3.99/|\lambda|^{0.34})}$ [^6] We assume that for $|\lambda|\gg 10^{3}$, $p$ and $\tilde{L}$ asymptomatically approach the ensemble average values derived for $|\lambda|=10^{3}$. with considerable scatter above and below this relation of $\sim 4$ depending on inclination angles and phase $\Delta\Phi$. Nevertheless, the trend is robust, with $\tilde{L}$ decreasing, reaching an asymptotic value for $|\lambda|\gg 1$ corresponding to an unmagnetized companion. Secondly, we also notice the correlation in Figure 19 for the average index, $\overline{p}$ (averaged over all the cases with equal $|\lambda|$), with $|\lambda|$. We find $\overline{p}\sim(6\pm 0.2)-(3.6\pm 0.1)|\lambda|^{-0.21\pm 0.02}$. Indeed, $\overline{p}$ is also correlated with $\tilde{L}$, thus $\overline{p}$ may be eliminated in favor of $\lambda$. Putting it all together yields a rough scaling law:
@@
×$\times\left(\frac{\Omega}{\Omega_{\rm i}}\right)^{(6\pm 0.2)-(3.6\pm 0.1)|\lambda|^{-0.21\pm 0.02}}\rm erg\,s^{-1}$ (15)
[…]
@@
Figure 21: A snapshot of the forces that develop at each star for the case with $\alpha_{1}=0^{\circ},\,\alpha_{2}=45^{\circ}$ and $\lambda=1$ (
@@ 3.2 Forces and Torques
In Section 3.1.2, we noted that when the magnetic moment of one of the two stars is not parallel to the orbital axis, the EM energy outflow becomes asymmetric between the north and the south orbital hemispheres. This behavior implies the existence of backreaction forces on the stars, generating torques that could eventually have an impact on the system’s orbital dynamics as well as individual NS spins.
@@
and the force density [^7] J. Pétri (2021) used the Laplace force to calculate the torque for a NS with an off-centered dipole magnetic moment. However, that approach can not be used here because flux evolves over time (V. S. Beskin & A. A. Zheltoukhov, 2014). Therefore in Eq. 17 the term involving $\partial\textbf{L}_{\rm Poynting}/\partial t$ is required. reads
@@
To further validate this approach, we applied the method to single, isolated pulsar cases. For a centered magnetic dipole moment, the net force was found to be zero, as expected. Moreover, the opposite forces exerted on the stellar surface produced a torque that led to a spin-down consistent with the corresponding EM spin-down power. For the case of an off-centered dipole moment, where the so-called rocket effect (E. R. Harrison & E. Tademaru, 1975; J. Pétri, 2021; V. Agalianou & K. N. Gourgouliatos, 2023) induces a net force, our calculations successfully reproduced the expected results, confirming the accuracy of our implementation.
In the BNS system, we find the EM forces can be several orders of magnitude smaller than gravitational forces, consistent with expectations of vacuum and magnetostatic calculations (e.g., M. Lira et al., 2022). Even for magnetar-like stars with strong magnetic fields ($\sim 10^{15}-10^{16}$ G), the EM forces still remain at least two to three orders of magnitude smaller than gravitational forces, being two orders of magnitude smaller only during the final moments before the merger. Yet, such ${\cal O}(1\%)$ effects are possibly discernible in third-generation gravitational wave detectors for loud mergers with waveform signal-to-noise exceeding $10^{3}$ (S. Borhanian & B. S. Sathyaprakash, 2024). Interestingly, in cases where the Poynting flux is asymmetric between the two hemispheres, the forces develop a component along the $z$ axis, perpendicular to the orbital plane (see Figure 21). These forces generate a torque along the $y$ axis in the co-rotating frame. This torque not only alternates in sign but also increases in magnitude over time, gradually tilting the orbital plane. Figure 22 schematically illustrates the increasing force components at their peak amplitudes, which drive the torque at two snapshots where the torque has an opposite sign. As a result, the stars exhibit oscillatory motion relative to an orbital plane that gradually tilts over time.
@@
The evolution of the torque and its effects are shown in Figure 23. The top row shows the torque along the $y$-axis for different magnetic inclination angles $\alpha_{2}$ ($0^{\circ}$, $45^{\circ}$, and $90^{\circ}$) and magnetic moment ratios $\lambda$ (both positive and negative values). The bottom row presents the corresponding evolution of the orbital plane deviation angle, $\theta$, induced by this torque. For stars with magnetic fields of $\sim 3\times 10^{15}$ G, the deviation angle increases up to $\sim 10^{-5}$ rad during the final $7.7$ ms of the system’s evolution. While this angle appears small, it demonstrates a clear trend of continuous growth over time. Considering that the total inspiral duration significantly exceeds $7.7$ ms, the cumulative effect of this torque from earlier times could potentially leave an observable imprint on the corresponding gravitational wave signal. However, it is crucial to highlight that these calculations were conducted in post-processing and did not account for the dynamical feedback of the magnetosphere on the orbital motion of the stars. For a more accurate and comprehensive assessment, the evolution should be treated self-consistently by dynamically incorporating this feedback into simulations, updating the stars’ positions at each time step. Investigating the full significance of this EM backreaction physics is reserved for future work.
@@ 3.2.1 Electromagnetic Crustal Deformations And Stresses
During the rapid evolution of the field in the magnetosphere, the forces and torques will result in nonuniform stresses on the NS crusts. This, if sufficiently large, could provoke a failure of the crustal Coulomb lattice. In order to determine whether the non-elastic response of the crust exceeds its yield limit, we follow the procedure described in S. K. Lander et al. (2015).
@@
Considering a maximum breaking strain $\kappa_{\rm\max}\sim 0.1$ (C. J. Horowitz & K. Kadau, 2009) the crustal lattice will yield (“crack”) if the von Mises criterion $K=\sqrt{\kappa_{ij}\kappa^{ij}/2}\geq\kappa_{\rm\max}$ is satisfied.
Using as an example the case with $\alpha_{1}=0^{\circ},\alpha_{2}=45^{\circ},\lambda=-1$ and $\Delta\Phi=0^{\circ}$ ( $\uparrow$ $\downarrow$ ) at $\tau\approx-30$ ms, we find that $K$ ranges from $\sim 0.2$ to $\sim 0.8\times 10^{-6}B_{\rm 12}^{2}$ for the zero-inclination star while reaching $\sim 1.2\times 10^{-6}B_{\rm 12}^{2}$ in the case of the inclined star. These values are non-uniform over the surfaces of the stars and vary with time with the orbital motion. The distribution of $K$ on the star surface could be seen for the case with $\alpha_{1}=0,\alpha_{2}=45^{\circ},\lambda=1$ and $\Delta\Phi=0$ ( $\uparrow$ $\uparrow$ ) in Figure 21. The stresses on the crust thus vary on a timescale $\sim\pi/\Omega$. At late times in the inspiral, the motion of stresses may approach or exceed the typical shear wave propagation speed in the crust, possibly exciting crustal normal modes (L. Samuelsson & N. Andersson, 2007).
These values suggest that if one of the NSs has a magnetic field of $\sim 10^{15}\rm,G$, the resulting EM stresses could induce significant deformations, potentially leading to catastrophic crustal failure. Such an event would likely produce an X-ray transient similar to a magnetar giant flare, while the excitation of crustal normal modes could give rise to additional observational signatures (e.g., E. R. Most et al., 2024).
@@ 3.3 High-energy Emission
In particular, studies of global pulsar magnetospheres, including FF (X.-N. Bai & A. Spitkovsky, 2010; I. Contopoulos & C. Kalapotharakos, 2010), dissipative (C. Kalapotharakos et al., 2012, 2014; J. Li et al., 2012; G. Cao & X. Yang, 2019), and kinetic plasma models (A. Y. Chen & A. M. Beloborodov, 2014; B. Cerutti et al., 2016; G. Brambilla et al., 2018; C. Kalapotharakos et al., 2018; A. A. Philippov & A. Spitkovsky, 2018; C. Kalapotharakos et al., 2023; B. Cerutti et al., 2024) have established that the primary region for particle acceleration is near the equatorial current sheet beyond the light cylinder. In this region, particles are accelerated by accelerating electric field components $\textbf{E}_{\rm acc}$, producing high-energy photons.
C. Kalapotharakos et al. (2019, 2022, 2023) demonstrated that the Fermi-LAT gamma-ray pulsars follow a “fundamental plane” gamma-ray luminosity relation ($L_{\rm\gamma}=10^{14.3\pm 1.3}\epsilon_{\rm cut}^{1.39\pm 0.17}B_{\star}^{0.12\pm 0.03}\dot{\cal{E}}^{0.39\pm 0.05}$ erg/s, where $\epsilon_{\rm cut}$, the spectral cutoff energy, is measured in MeV, $B_{\star}$, the surface magnetic field, in G, and $\dot{\cal{E}}$, the spin-down power, in erg/s). Remarkably, this correlation of observables aligns closely with the theoretical relations predicted for curvature radiation ($L_{\rm\gamma}\propto\epsilon_{\rm cut}^{4/3}B_{\star}^{1/6}\dot{\mathcal{E}}^{5/12}$). Thus, curvature radiation is adopted as the dominant emission mechanism in this work.
@@ 3.3.1 Maximum Particle Energies
The Lorentz factor, $\gamma_{L}$, is determined by the balance between energy gain due to acceleration by the $E_{\rm acc}$ and energy loss from curvature radiation. This balance (in the ultra-relativistic limit) is governed by the equation
$\frac{d\gamma_{L}}{dt}=\frac{q_{e}E_{\rm acc}}{m_{e}c}-\frac{2q_{e}\gamma_{L}^{4}}{3R_{C}^{2}m_{e}c},$ (19)
@@
$\gamma_{L}=\left(\frac{3R_{C}^{2}E_{\rm acc}}{2}\right)^{1/4}.$ (20)
For the double NS system, we now conjecture that particles have reached the curvature radiation-reaction limit. Furthermore, we consider that the radius of curvature, $R_{C}$, is the same order of magnitude as the radius of the orbital lengthscale $R_{C}\sim R_{\rm LC}$, and that $E_{\parallel}$ is assumed to be comparable to the magnetic field at the light cylinder, i.e., $E_{\rm acc}\sim B_{\rm LC}\approx B_{\star}(r_{\star}/R_{\rm LC})^{3}$. Then $\gamma_{L}$ can be derived using Eq. (20). Substituting this value into Eq. (18) yields the characteristic photon energy of the curvature radiation spectrum.
@@
During the last few milliseconds before coalescence, the putative cutoff energy attains $\sim 10$ TeV. For systems with stronger magnetic fields, the implied cutoff energy can be even higher. As highlighted in the right panel of Fig. 24, even for stars with $10^{10}$ G, the cutoff energy might attain $\sim 1$ TeV just $1$ ms prior to the merger, while for stars with $10^{15}$ G it may exceed 1 PeV. However, this is not the full story as gamma-ray opacities to electron-positron pair production (single photon or two photons) rise drastically close to coalescence, either inhibiting escape of photons of these energies or screening $E_{\rm acc}$ (see §3.3.3).
@@ 3.3.2 Gamma-ray Emission Anisotropy and Light Curves
In binary systems, however, the situation becomes significantly more complex. The solid-body rotation assumption no longer holds due to the orbital motion and the continuous reconfiguration of the magnetic field structure caused by the motion of the irrotational stars. As a result, the field evolves dynamically on timescales comparable to the orbital period. Additionally, identifying the stellar surface regions where high-energy particles originate at each time step is non-trivial. In the isolated pulsar case, particles in the current sheet can reliably be traced to the polar cap rim, but in binary systems, determining these regions does not guarantee that the particles will reach the dissipative regions of the magnetosphere. A proper treatment would require integrating particle trajectories originating from all points on the stellar surfaces at every time step, which is computationally expensive given the rapid evolution of the magnetic field.
@@
More specifically, we calculate $E_{\rm acc}$ in a dense grid of spherical shells ($dr_{\rm shell}=0.1r_{\star}$) spanning from $0.9R_{\rm LC}$ to $2R_{\rm LC}$. Since $R_{\rm LC}$ decreases as the orbital frequency increases, the volume of the shell evolves over time. In FF solutions, $E_{\rm acc}$ is negligible, so we follow the prescription of C. Kalapotharakos et al. (2014) to derive $E_{\rm acc}$, i.e., the parallel (to the magnetic field) electric field component. In this framework, $E_{\rm acc}$ is expressed as
$\frac{d\textbf{E}_{acc}}{dt}=c\mathbf{\nabla}\times\textbf{B}+\frac{4\pi}{\sigma}\textbf{J}_{||}.$ (22)
@@
To avoid inadvertently selecting regions where (×$(\nabla\times\mathbf{B})_{\parallel}$ appears high simply because they are located in proximity to strong magnetic fields (e.g., near the stellar surfaces), we apply additional constraints to ensure the identified accelerating regions are physically meaningful. Specifically, we identify locations where $E{\rm{}_{acc}}$ exceeds 10% (to avoid noise in the simulation) of the local electric field magnitude, $E$, as the regions where high-energy emission originates. Figure 25 illustrates these regions at two-time snapshots, separated by half an orbital period, for $\alpha_{1}=0^{\circ}$, $\alpha_{2}=45^{\circ}$, $\Delta\Phi=0^{\circ}$ and $\lambda=-1$ ( $\uparrow$ $\downarrow$ ). These regions are predominantly located near the current sheets, consistent with single pulsar expectations, and they evolve dynamically as the magnetosphere reconfigures during the system’s inspiral.
@@
A major challenge in approaches involving test particle trajectories lies in determining how many physical particles are represented by each test particle. This uncertainty, which can vary from one test particle to another, persists even in single pulsar studies and can only be partially resolved with more self-consistent PIC simulations. To mitigate this issue for binary systems, we adopt a simpler prescription where the emitting points lie on spherical surfaces corresponding to a specific Poynting flux value. The relative weights of these emission points are determined by the strength of their local $E_{\rm acc}$ components. To include the high-energy efficiency of each surface, we scale the total emissivity relative to the Poynting flux power crossing that surface. Note here that in order to preserve causality, we use the flux that crosses the spherical surface containing the point where emission occurs at the time of the emission.
Figure 26: The evolution of high energy emission skymaps for BNS systems with $\alpha_{1}=0$, $\alpha_{2}=45^{o}$, $\Delta\Phi=0$ and $\lambda=1$(
@@
Algorithmically, at every time step, we divide the region between $0.9R_{\rm LC}$ and $2R_{\rm LC}$ into spherical shells centered on the system’s center. For each shell, we calculate the Poynting flux at its mid-radius surface and identify the emission points where $E_{\rm acc}/E$ exceeds the defined threshold. The emission from each sampled point is then weighted proportionally to $\sqrt{L_{\rm shell}}/\sum_{N_{\rm shell}}E_{\rm acc}$, where $L_{\rm shell}$ is the Poynting flux for the shell and $N_{\rm shell}$ is the number of sampled emission points within the shell.
Unlike the isolated pulsar case, where the nearly constant spin period allows for phase folding and coherent photon contributions, the rapidly evolving orbital motion in the binary system prevents the existence of a well-defined or coherent phase. Instead, the relative detection time for each photon must be calculated at a distant surface.
Figure 27: In the first row the high energy emission skymaps for $\alpha_{1}=45^{o}$, $\alpha_{2}=90^{o}$, $\Delta\Phi=0$ and $\lambda=1$ (
@@
where $r$ is the position vector of the emission point and $\boldsymbol{v}$ is the particle velocity. Since we assume equal masses, the center of mass coincides with the origin of the coordinate system. The relative detection time, measured with respect to a fiducial photon emitted from the system’s center of mass at the beginning of the simulation, at which each photon contributes to the observed emission, is then given by
@@
Figure 28: Optical depth for the case of $\alpha_{1}=0,\alpha_{2}=45^{\circ},\lambda=-1$ and $\Delta\Phi=0$ (
@@
) with $10^{12}$ G (Left) and $10^{14}$ G (Right) magnetic field stars. A wider line indicates a later stage in the evolution. The vertical dashed line highlights the minimum pair production threshold of $1$ MeV while the blue horizontal line is the ${\cal T}=1$ line. Note that in the right-hand panel, $\mathcal{T}$ may be overestimated near $\epsilon=1$ MeV (see main text for discussion), potentially permitting the escape of marginally higher-energy photons.
@@
In those figures, the vertical and horizontal axes denote the polar $\theta$ and azimuthal $\phi$ angles on a distant sphere around the system. The color indicates the high energy flux density that is beamed towards the direction defined by the angles $\theta$ and $\phi$ at the indicated time.

### v2

@@ 3.1.5 Scaling Laws for Poynting Luminosity and the Binary Braking Index
Empirically fitting the gross trend from our simulations $\langle L\rangle\approx~\tilde{L}\Omega^{p}$, we find an expected correlation between $\lambda$ and the overall normalization $\tilde{L}$ of the Poynting luminosity. Namely, $\tilde{L}/(10^{42}\,B_{12}^{2}r_{12\,\rm km}^{2}\rm\,erg\,s^{-1})~\sim~0.8\exp{(3.99/|\lambda|^{0.34})}$ [^7] We assume that, for $|\lambda|\gg 10^{3}$, $p$ and $\tilde{L}$ asymptomatically approach the ensemble average values derived for $|\lambda|=10^{3}$. with considerable scatter above and below this relation of $\sim 4$ depending on inclination angles and phase $\Delta\Phi$. Nevertheless, the trend is robust, with $\tilde{L}$ decreasing, reaching an asymptotic value for $|\lambda|\gg 1$ corresponding to an unmagnetized companion. Second, we also notice the correlation in Figure 19 for the average index, $\overline{p}$ (averaged over all the cases with equal $|\lambda|$), with $|\lambda|$. We find $\overline{p}\sim(6\pm 0.2)-(3.6\pm 0.1)|\lambda|^{-0.21\pm 0.02}$. Indeed, $\overline{p}$ is also correlated with $\tilde{L}$; thus, $\overline{p}$ may be eliminated in favor of $\lambda$. Putting it all together yields a rough scaling law:
@@
$\times\left(\frac{\Omega}{\Omega_{\rm i}}\right)^{(6\pm 0.2)-(3.6\pm 0.1)|\lambda|^{-0.21\pm 0.02}}\hskip-17.07164pt\rm erg\,s^{-1}$ (15)
where the $\langle\cdot\cdot\cdot\rangle$ denotes an orbital average, implicitly encoding the pulsations in Figure 12, while $\overline{\cdot\cdot\cdot}$ denotes the average over various combinations of $\alpha_{1}$, $\alpha_{2}$, $\Delta\Phi$, and the sign of $\lambda$ (see the dotted lines in Figures 12 and 20). Thus, Eq. (15) primarily captures the dependence of $\overline{\langle L\rangle}$ on $\Omega$ as function of $|\lambda|$.
@@
Figure 21: A snapshot of the forces that develop at each star for the case with $\alpha_{1}=0^{\circ},\,\alpha_{2}=45^{\circ}$, and $\lambda=1$ (
@@ 3.2 Forces and Torques
In Section 3.1.2, we noted that, when the magnetic moment of one of the two stars is not parallel to the orbital axis, the EM energy outflow becomes asymmetric between the north and the south orbital hemispheres. This behavior implies the existence of backreaction forces on the stars, generating torques that could eventually have an impact on the system’s orbital dynamics as well as individual NS spins.
@@
and the force density [^8] J. Pétri (2021) used the Laplace force to calculate the torque for an NS with an off-centered dipole magnetic moment. However, that approach cannot be used here because flux evolves over time (V. S. Beskin & A. A. Zheltoukhov, 2014). Therefore, in Eq. 17, the term involving $\partial\textbf{L}_{\rm Poynting}/\partial t$ is required. reads
@@
To further validate this approach, we applied the method to single, isolated pulsar cases. For a centered magnetic dipole moment, the net force was found to be zero, as expected. Moreover, the opposite forces exerted on the stellar surface produced a torque that led to a spin-down consistent with the corresponding EM spin-down power. For the case of an off-centered dipole moment in vacuum, where the so-called rocket effect (E. R. Harrison & E. Tademaru, 1975; J. Pétri, 2021; V. Agalianou & K. N. Gourgouliatos, 2023) induces a net force, our calculations successfully reproduced the expected results, confirming the accuracy of our implementation.
In the BNS system, we find the EM forces can be several orders of magnitude smaller than gravitational forces, consistent with expectations of vacuum and magnetostatic calculations (e.g., K. Ioka & K. Taniguchi, 2000; M. Lira et al., 2022). Even for magnetar-like stars with strong magnetic fields ($\sim 10^{15}-10^{16}$ G), the EM forces still remain at least 2 to 3 orders of magnitude smaller than gravitational forces, being 2 orders of magnitude smaller only during the final moments before the merger. Yet, such ${\cal O}(1\%)$ effects are possibly discernible in third-generation GW detectors for loud mergers with waveform signal-to-noise exceeding $10^{3}$ (S. Borhanian & B. S. Sathyaprakash, 2024). Interestingly, in cases where the Poynting flux is asymmetric between the two hemispheres, the forces develop a component along the $z$-axis, perpendicular to the orbital plane (see Figure 21). These forces generate a torque along the $y$-axis in the corotating frame. This torque not only alternates in sign but also increases in magnitude over time, gradually tilting the orbital plane. Figure 22 schematically illustrates the increasing force components at their peak amplitudes, which drive the torque at two snapshots where the torque has an opposite sign. As a result, the stars exhibit oscillatory motion relative to an orbital plane that gradually tilts over time.
@@
The evolution of the torque and its effects are shown in Figure 23. The top row shows the torque along the $y$-axis for different magnetic inclination angles $\alpha_{2}$ ($0^{\circ}$, $45^{\circ}$, and $90^{\circ}$) and magnetic moment ratios $\lambda$ (both positive and negative values). The bottom row presents the corresponding evolution of the orbital plane deviation angle, $\theta$, induced by this torque. For stars with magnetic fields of $\sim 3\times 10^{15}$ G, the deviation angle increases up to $\sim 10^{-5}$ rad during the final $7.7$ ms of the system’s evolution. While this angle appears small, it demonstrates a clear trend of continuous growth over time. Considering that the total inspiral duration significantly exceeds $7.7$ ms, the cumulative effect of this torque from earlier times could potentially leave an observable imprint on the corresponding GW signal. However, it is crucial to highlight that these calculations were conducted in postprocessing and did not account for the dynamical feedback of the magnetosphere on the orbital motion of the stars. For a more accurate and comprehensive assessment, the evolution should be treated self-consistently by dynamically incorporating this feedback into simulations, updating the stars’ positions at each time step. Investigating the full significance of this EM backreaction physics is reserved for future work.
@@ 3.2.1 Electromagnetic Crustal Deformations and Stresses
During the rapid evolution of the field in the magnetosphere, the forces and torques will result in nonuniform stresses on the NS crusts. This, if sufficiently large, could provoke a failure of the crustal Coulomb lattice. In order to determine whether the nonelastic response of the crust exceeds its yield limit, we follow the procedure described in S. K. Lander et al. (2015).
@@
Considering a maximum breaking strain $\kappa_{\rm\max}\sim 0.1$ (C. J. Horowitz & K. Kadau, 2009), the crustal lattice will yield (“crack”) if the von Mises criterion $K=\sqrt{\kappa_{ij}\kappa^{ij}/2}\geq\kappa_{\rm\max}$ is satisfied.
Using as an example the case with $\alpha_{1}=0^{\circ},\alpha_{2}=45^{\circ},\lambda=-1$, and $\Delta\Phi=0^{\circ}$ ( $\uparrow$ $\downarrow$ ) at $\tau\approx-30$ ms, we find that $K$ ranges from $\sim 0.2$ to $\sim 0.8\times 10^{-6}B_{\rm 12}^{2}$ for the zero-inclination star while reaching $\sim 1.2\times 10^{-6}B_{\rm 12}^{2}$ in the case of the inclined star. These values are nonuniform over the surfaces of the stars and vary with time with the orbital motion. The distribution of $K$ on the star surface could be seen for the case with $\alpha_{1}=0,\alpha_{2}=45^{\circ},\lambda=1$, and $\Delta\Phi=0$ ( $\uparrow$ $\uparrow$ ) in Figure 21. The stresses on the crust thus vary on a timescale $\sim\pi/\Omega$. At late times in the inspiral, the motion of stresses may approach or exceed the typical shear wave propagation speed in the crust, possibly exciting crustal normal modes (L. Samuelsson & N. Andersson, 2007).
These values suggest that if one of the NSs has a magnetic field of $\sim 10^{15}\rm G$, the resulting EM stresses could induce significant deformations, potentially leading to catastrophic crustal failure. Such an event would likely produce an X-ray transient similar to a magnetar giant flare, while the excitation of crustal normal modes could give rise to additional observational signatures (e.g., E. R. Most et al., 2024).
@@ 3.3 High-energy Emission
In particular, studies of global pulsar magnetospheres, including FF (X.-N. Bai & A. Spitkovsky, 2010; I. Contopoulos & C. Kalapotharakos, 2010), dissipative (C. Kalapotharakos et al., 2012, 2014; J. Li et al., 2012; G. Cao & X. Yang, 2019), and kinetic plasma models (A. Y. Chen & A. M. Beloborodov, 2014; B. Cerutti et al., 2016; G. Brambilla et al., 2018; C. Kalapotharakos et al., 2018; A. A. Philippov & A. Spitkovsky, 2018; C. Kalapotharakos et al., 2023; B. Cerutti et al., 2025), have established that the primary region for particle acceleration is near the equatorial current sheet beyond the light cylinder. In this region, particles are accelerated by accelerating electric field components $\textbf{E}_{\rm acc}$, producing high-energy photons.
[…]
@@ 3.3.1 Maximum Particle Energies
The Lorentz factor, $\gamma_{L}$, is determined by the balance between energy gain due to acceleration by the $E_{\rm acc}$ and energy loss from curvature radiation. This balance (in the ultrarelativistic limit) is governed by the equation
$\frac{d\gamma_{L}}{dt}=\frac{q_{e}E_{\rm acc}}{m_{e}c}-\frac{2q_{e}^{2}\gamma_{L}^{4}}{3R_{C}^{2}m_{e}c},$ (19)
@@
$\gamma_{L}=\left(\frac{3R_{C}^{2}E_{\rm acc}}{2q_{e}}\right)^{1/4}.$ (20)
For the double NS system, we now conjecture that particles have reached the curvature radiation-reaction limit. Furthermore, we consider that the radius of curvature, $R_{C}$, is the same order of magnitude as the radius of the orbital lengthscale $R_{C}\sim R_{\rm LC}$, and that $E_{\parallel}$ is assumed to be comparable to the magnetic field at the light cylinder, i.e., $E_{\rm acc}\sim B_{\rm LC}\approx B_{\star}(r_{\star}/R_{\rm LC})^{3}$. Then, $\gamma_{L}$ can be derived using Eq. (20). Substituting this value into Eq. (18) yields the characteristic photon energy of the curvature radiation spectrum.
@@
During the last few milliseconds before coalescence, the putative cutoff energy attains $\sim 10$ TeV. For systems with stronger magnetic fields, the implied cutoff energy can be even higher. As highlighted in the right panel of Fig. 24, even for stars with $10^{10}$ G, the cutoff energy might attain $\sim 1$ TeV just $1$ ms prior to the merger, while for stars with $10^{15}$ G it may exceed 1 PeV. However, this is not the full story as gamma-ray opacities to electron-positron pair production (single photon or two photons) rise drastically close to coalescence, either inhibiting the escape of photons of these energies or screening $E_{\rm acc}$ (see §3.3.3).
@@ 3.3.2 Gamma-ray Emission Anisotropy and Light Curves
With the assumption of synchronized NSs, where the stars have the same orbital and spin frequencies, the magnetosphere structure corotates with the system, and thus, the above approaches from single pulsar studies are still applicable. For such a case, in N. Ortiz et al. (2022), the authors considered a spherical conducting surface that encloses the system and identifying there the polar caps and the “separatrix”, and they construct high-energy skymaps and light curves following the prescription in X.-N. Bai & A. Spitkovsky (2010).
In more generic binary systems, however, the situation becomes significantly more complex. The solid-body rotation assumption no longer holds due to the orbital motion and the continuous reconfiguration of the magnetic field structure caused by the motion of the irrotational stars. As a result, the field evolves dynamically on timescales comparable to the orbital period. Additionally, identifying the stellar surface regions where high-energy particles originate at each time step is nontrivial. In the isolated pulsar case, particles in the current sheet can reliably be traced to the polar cap rim, but in binary systems, determining these regions does not guarantee that the particles will reach the dissipative regions of the magnetosphere. A proper treatment would require integrating particle trajectories originating from all points on the stellar surfaces at every time step, which is computationally expensive given the rapid evolution of the magnetic field.
@@
More specifically, we calculate $E_{\rm acc}$ in a dense grid of spherical shells ($dr_{\rm shell}=0.1r_{\star}$) spanning from $0.9R_{\rm LC}$ to $2R_{\rm LC}$. Since $R_{\rm LC}$ decreases as the orbital frequency increases, the volume of the shell evolves over time. In FF solutions, $E_{\rm acc}$ is negligible, so we follow the prescription of C. Kalapotharakos et al. (2014) to derive $E_{\rm acc}$, i.e., the parallel (to the magnetic field) electric field component. In this framework, $E_{\rm acc}$ is derived from
$\frac{d\textbf{E}_{acc}}{dt}=c\mathbf{(}{\nabla}\times\textbf{B})_{||}-4\pi\textbf{J}_{||}$ (22)
and expressed as
$\textbf{E}_{acc}=\frac{c\mathbf{(}{\nabla}\times\textbf{B})_{||}}{4\pi\sigma}.$ (23)
@@
To avoid inadvertently selecting regions where $(\nabla\times\mathbf{B})_{\parallel}$ appears high simply because they are located in proximity to strong magnetic fields (e.g., near the stellar surfaces), we apply additional constraints to ensure the identified accelerating regions are physically meaningful. Specifically, we identify locations where $E{\rm{}_{acc}}$ exceeds 10% (to avoid noise in the simulation) of the local electric field magnitude, $E$, as the regions where high-energy emission originates. Figure 25 illustrates these regions at two-time snapshots, separated by half an orbital period, for $\alpha_{1}=0^{\circ}$, $\alpha_{2}=45^{\circ}$, $\Delta\Phi=0^{\circ}$, and $\lambda=-1$ ( $\uparrow$ $\downarrow$ ). These regions are predominantly located near the current sheets, consistent with single pulsar expectations, and they evolve dynamically as the magnetosphere reconfigures during the system’s inspiral.
@@
A major challenge in approaches involving test particle trajectories lies in determining how many physical particles are represented by each test particle. This uncertainty, which can vary from one test particle to another, persists even in single pulsar studies and can only be partially resolved with more self-consistent PIC simulations. To mitigate this issue for binary systems, we adopt a simpler prescription where the emitting points lie on spherical surfaces corresponding to a specific Poynting flux value. The relative weights of these emission points are determined by the strength of their local $E_{\rm acc}$ components. To include the high-energy efficiency of each surface, we scale the total emissivity relative to the Poynting flux power crossing that surface. Note here that, in order to preserve causality, we use the flux that crosses the spherical surface containing the point where emission occurs at the time of the emission.
Figure 26: The evolution of high-energy emission skymaps for BNS systems with $\alpha_{1}=0$, $\alpha_{2}=45^{o}$, $\Delta\Phi=0$, and $\lambda=1$(
@@
Algorithmically, at every time step, we divide the region between $0.9R_{\rm LC}$ and $2R_{\rm LC}$ into spherical shells centered on the system’s center. For each shell, we calculate the Poynting flux at its midradius surface and identify the emission points where $E_{\rm acc}/E$ exceeds the defined threshold. The emission from each sampled point is then weighted proportionally to $\sqrt{L_{\rm shell}}/\sum_{N_{\rm shell}}E_{\rm acc}$, where $L_{\rm shell}$ is the Poynting flux for the shell, and $N_{\rm shell}$ is the number of sampled emission points within the shell.
Unlike the isolated pulsar case, where the nearly constant spin period allows for phase folding and coherent photon contributions, the rapidly evolving orbital motion in the binary system prevents the existence of a well-defined or coherent phase (N. Ortiz et al., 2022). Instead, the relative detection time for each photon must be calculated at a distant surface.
Figure 27: In the first row the high-energy emission skymaps for $\alpha_{1}=45^{o}$, $\alpha_{2}=90^{o}$, $\Delta\Phi=0$, and $\lambda=1$ (
@@
where $r$ is the position vector of the emission point, and $\boldsymbol{v}$ is the particle velocity. Since we assume equal masses, the center of mass coincides with the origin of the coordinate system. The relative detection time, measured with respect to a fiducial photon emitted from the system’s center of mass at the beginning of the simulation, at which each photon contributes to the observed emission, is then given by
@@
Figure 28: Optical depth for the case of $\alpha_{1}=0,\alpha_{2}=45^{\circ},\lambda=-1$, and $\Delta\Phi=0$ (
@@
) with $10^{12}$ G (left) and $10^{14}$ G (right) magnetic field stars. A wider line indicates a later stage in the evolution. The vertical dashed line highlights the minimum pair-production threshold of 11 MeV while the blue horizontal line is the ${\cal T}=1$ line. Note that, in the right-hand panel, $\mathcal{T}$ may be overestimated near $\epsilon=1$ MeV (see main text for discussion), potentially permitting the escape of marginally higher-energy photons.
@@
In those figures, the vertical and horizontal axes denote the polar $\theta$ and azimuthal $\phi$ angles on a distant sphere around the system. The color indicates the high-energy flux density that is beamed toward the direction defined by the angles $\theta$ and $\phi$ at the indicated time.